OpenStudy (anonymous):

MEDAL AND FAN For the following system, use the second equation to make a substitution for y in the first equation. 3 x + y = 2 y + 4 = 3 x ANSWERS: 3 x + 3 x - 4 = 2 3 x + 3 - 4 x = 2 3 x + 3 x + 4 = 2

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 3 x + y = 2 }\) \(\large\color{blue}{ \rm y + 4 = 3 x }\) \(\large\color{blue}{ \rm 3 x + y = 2 }\) \(\large\color{blue}{ \rm 3x-y = 4 }\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i dont understand

hartnn (hartnn):

y + 4 = 3 x since 4 is ADDED to 'y' SUBTRACT 4 from both sides to isolate 'y' what u get ?

OpenStudy (anonymous):

May I ask what type of equation is this called? I remember doing this ages ago...I want to learn this again.

OpenStudy (anonymous):

im getting so confused by this subject. its hard for me to grasp this

hartnn (hartnn):

3 x + y = 2 y + 4 = 3 x is a system of linear equations with 2 linear equation. to solve them simultaneously, (to find x and y) there are several methods, i) elimination, ii) substitution and some others... what solmon suggested is elimination method what question expects and i suggested is substitution method

OpenStudy (anonymous):

You're not alone spasekat, hahaha. I'm confused here too. Thanks hartnn! :)

hartnn (hartnn):

y + 4 = 3 x subtract 4 from both sides can you do that ?

OpenStudy (anonymous):

do i subtract 4 from x?

hartnn (hartnn):

right side = 3x if you subtract 4, it'll be 3x-4 they won't combine , because they are NOT like terms

hartnn (hartnn):

what about left side ?

OpenStudy (anonymous):

y - 4?

OpenStudy (anonymous):

or y + (-4)

hartnn (hartnn):

it was y+4 you subtract 4 y+4-4 what does that equal ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

y + 4-4 = 3x - 4

hartnn (hartnn):

yes 4-4 =0 and y+0 = y so, we have \(\Large y = 3x-4\) clear ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

now look at 1st equation 3x+y = 2 but we already got y = 3x-4 so replace 'y' in first equation with '3x-4' what do u get ?

OpenStudy (anonymous):

3x + 3x - 4 = 2

hartnn (hartnn):

good! thats your first option :) any doubts ?

OpenStudy (anonymous):

no :D

OpenStudy (anonymous):

that was really helpful! i am a fan and medal now for you

hartnn (hartnn):

thanks :D